The number of photons of light of ν reciprocal =2.5*10^6 m^-1 necessary to provide 1J of energy are
a) 2*10^18.                   b)2*10^17
c)2*10^20.                     d) 2*10^19
         

Grade:11

1 Answers

Jobin C J
5 years ago
E=nhν
 
But ν=c/λ
 
So we can write E=nhc/λ
 
E=nhc/λ  ⇒  E=nhc×1/λ
 
But ṽ=1/λ
 
So E=nhcṽ
 
E=1J,  h=6.63×10-34Js, C=3×108ms-1, ṽ=2.5×106m-1
 
n= E/(hcṽ)
 
n= 1/(6.63×10-34 × 3×108 × 2.5×106)
 
n= 2×1018
 
Option 'a' is correct.